OpenStudy (anonymous):

please help me :/ What is the quotient x^2-16 /2x^2-9x+4 divided by 2x^2+14x+24/4x+4 ? Choices : 1/x+3 2x+2/x+3 2x+2/2x^2-5x-3 2(x+1)/2x^2-5x-3

OpenStudy (anonymous):

mabye this ... What is the quotient ((x^2-16)/(2x^2-9x+4))/((2x^2+14x+ 24)/(4x+4)) 1. x^2-16 = (x-4)(x+4) 2. 2x^2-9x+4 = (2x-1)(x-4) 3. (2x^2+ 14x + 24) = 2(x^2+7+12) = 2 (x+3)(x+4) 4. (4x+4) = (4(x+1))

OpenStudy (anonymous):

so was that right?